let f(x)=-2+3 and g (x)=x2 (x squared)
f(g(3))

Answer by sdmmadam@yahoo.com(530)

let f(x)=-2+3 and g (x)=x2 (x squared)
I think f(x) is typed wrong. Either it should be f(x)=-2x+3 or f(x) =-2+3x
1) let f(x) = -2x+3 ----(1)
and g(x) = x^2 ----(2)
f(g(3))= f[3^2] using (2) (putting x=3 in (2))
=f(9)
=-2X9+3 = -18+3 = -15
Answer: f(g(3))=-15
2) let f(x) = -2+3x ----(1)
and g(x) = x^2 ----(2)
f(g(3))= f[3^2] using (2) (putting x=3 in (2))
=f(9)
=-2+3X9 = -2+27 = 25
Answer: f(g(3))= 25